Sox Engineering designs and constructs air conditioning and heating​ (HVAC) systems for hospitals and clinics. Currently the​ company's staff is overloaded with design work. There is a major design project due in

99

weeks. The penalty for completing the design late is

​$14 comma 00014,000

per​ week, since any delay will cause the facility to open later than​ anticipated, and cost the client significant revenue. If the company uses its inside engineers to complete the​ design, it will have to pay them overtime for all work. Sox has estimated that it will cost

​$11 comma 00011,000

per week​ (wages and​ overhead), including late​ weeks, to have company engineers complete the design. Sox is also considering having an outside engineering firm do the design. A bid of

​$90 comma 00090,000

has been received for the completed design. Yet another option for completing the design is to conduct a joint design by having a third engineering company complete all electromechanical components of the design at a cost of

​$60 comma 00060,000.

Sox would then complete the rest of the design and control systems at an estimated cost of

​$32 comma 00032,000.

Sox has estimated the following probabilities of completing the project within various time frames when using each of the three options. Those estimates are shown in the following​ table:

Probability of Completing the Design

Option

On Time

One Week Late

Two Weeks Late

Three Weeks Late

Internal Engineers

0.30.3

0.40.4

0.30.3

long dash—

External Engineers

0.10.1

0.40.4

0.30.3

0.20.2

Joint Design

0.20.2

0.30.3

0.40.4

0.10.1

The expected monetary value using inside engineers to complete the design is

?????

​(Enter

your response as an

integer.​)

The expected monetary value using an outside engineering firm to complete the design is ???

​(Enter

your response as an

integer.​)

The expected monetary value of conducting a joint design is ??

​(Enter

your response as an

integer.​)

What is the best decision based on an expected monetary value​ criterion?

Joint designJoint design

External engineering firmExternal engineering firm

Internal engineers
Answer #1

The penalty for completing design late is $14000 per week. This is common for all the decisions.

The internal design will incur an additional overtime cost of $11000 per week. This makes it $25000 per week effective cost. However, there are no other cost associated with this.

The external design will incur a one-time cost of $90000. In addition if the design is late, the late fee of $14000 is applicable.

The join design will cost $60000 + $32000 = $92000. In addition, if the design is late, there will be the delaying cost of $14000 per week.

Expected monetary value using inside engineers = 0.3*0 + 0.4*25000 + 0.3*25000*2 = 25000

Expected monetary value using external engineers = 0.1*90000 + 0.4*(90000+14000) + 0.3*(90000+14000*2)+0.2*(90000+14000*3) = 112400

Expected monetary value using joint design =0.2*92000 + 0.3*(92000+14000) + 0.4*(92000+14000*2)+0.1*(92000+14000*3) = 111600

According the EMV criterion, the best design is inside engineers (as this is the cheapest option)